OpenStudy (anonymous):

What is the slope of a line that passes through the point (−2, 3) and is parallel to a line that passes through (3, 7) and (−2, −8)?

OpenStudy (anonymous):

@Directrix

OpenStudy (anonymous):

Hi HelloGoodmorning Recommended you :)

Directrix (directrix):

What is the slope of the line determined by these two points: (3, 7) and (−2, −8)

OpenStudy (anonymous):

The y is the first number right? like -2 and 3

OpenStudy (anonymous):

Sorry that was stupid to ask I see it in the formula now

Directrix (directrix):

I don't think of it that way.

Directrix (directrix):

(-8 - 7)/(-2 - 3) = ? for the slope of that line?

OpenStudy (anonymous):

15/-5

Directrix (directrix):

I did not get that. (-8 - 7)/(-2 - 3) = -15/(-5) = ? Negative divided by negative is ?

OpenStudy (anonymous):

Positive

Directrix (directrix):

So the slope is +3

OpenStudy (anonymous):

3

OpenStudy (anonymous):

So thats it?

Directrix (directrix):

Look at the posted question: What is the slope of a line that passes ... Did you mean to write this --> What is the *equation* of a line that passes .... Check the problem, please.

OpenStudy (anonymous):

yeo

OpenStudy (anonymous):

i already asked that @Directrix

OpenStudy (anonymous):

the slope is right

OpenStudy (anonymous):

Yeah It's aasking for the slope

Directrix (directrix):

I find the inclusion of this information odd or superfluous: *** that passes through the point (−2, 3)***

OpenStudy (anonymous):

Lol dude for some reason that made me laugh

Directrix (directrix):

Me, too. @SOS101 Are the answer options equations or just numbers?

OpenStudy (anonymous):

Lol sadly its a fill in the blank

Directrix (directrix):

@JOELMATHEWS1234 >>i already asked that If you have already done this problem, why are we doing it again?

OpenStudy (anonymous):

no , asked wether it is asking for slope or equation :P

Directrix (directrix):

@SOS101 Do this: Check the problem source and make sure you posted the problem correctly. @JOELMATHEWS1234 @SOS101 needs to check is posted here against the problem source. If this is the problem as posted, then the answer is 3. Otherwise ....

OpenStudy (anonymous):

The problem sourse is asking for the slope

OpenStudy (anonymous):

Thanks @Directrix

Directrix (directrix):

It is over, then. 3 is the slope.

Directrix (directrix):

You are welcome.

OpenStudy (anonymous):

:)